Forces in a Three-Charge System 5 of 14 ConstantsI Periodic Table Part A Coulombs law for the magnitude of the force F between two particles with charges Q and Q separated by a distance d is What is the net force exerted by these two charges on a third charge 52.0 nC placed between q and g2 at Zg=-1.160 m? Your answer may be positive or negative, depending on the direction of the force Express your answer numerically in newtons to three significant figures. QQ1 where K =-, and 4neo View Available Hint(s) 8.854 x 1012 C2 /(N m2) is the permittivity of free space Consider two point charges located on the x axis one charge, q1 17.5 nC, is located at 11.650 m the second charge, 92-31.0 nC is at the origin (z = 0.0000) Force on 3.11.10 Submit Incorrect: Try Again: 2 attempts remainingPeden 5) Consider two points charや&add ση rhe aC-axis s 52010 9.45 A132 rede 汐 .109 15 31 0 2.03-I 732 1.08.10 3.11-10-s =







Hello, I am not sure what part I miss.

Do you what to fix to get the correct answer?

Since there is only two chances remained, it would be very helpful if you provide me with the answer.

Thank you.

***Answer is neither -3.11*10^-5 nor 3.11*10^-5.
Answer #1

1.650m q1 -17.5nC q3 52.0nC q2-31.0nC 0.49m F1 F2

Like charges repel and unlike charges attracts

so q3 will be moving towards q1 and away from q2

F=KQQ'/d^{2}

F_{1}=KQ_{1}Q_{3}/d^{2}

F, = 8.99 * 109 * 17.5 * 10-9C * 52 * 10-9C)(0.49m)2

(In written answer, mistake is made on values of F31. Red is the correction )

F1 = 3.41 * 10-5N(towards negative side)

=====================

F_{2}=KQ_{2}Q_{3}/d^{2}

F2 8.99 109 311052 10C/(1.160m)-

F21.077 10(towards negative side)

====================

F3 = F1 + F2

F3 3.41 101.077 10 N

F3 = 4.487 * 10-5M(towards negative side)

ANSWER: F_{3}=-4.487*10^{-5}N

==============================